Date of Lab Numbers 1-4: Apply rule (a) or (b). Show every step of your work, including units. I. 41.57s + 0.016s 2. (1.57m)(12.02m)- 3. 1502 kg-3.14 kg Apply rules (a) and (b) at every step, showing work. Remember, order of operations requires division before addition. Do not find common denominators with this problem. 4. 999.0m+888m 444m- 6.00s 5.00s 4.80s work #4 again, by the simplified procedure (Rule C) for calculations with measured data. Show only this much work: The original problem, the calculator result, and your final (rounded) answer. 6. Find the percent difference between 0.058 g and 0.043 g, with respect to the 0.043 g value.
7. The known value is 980 cm/s, but a student measures 1010 cm/s. Find the percent error. 8. Regarding time as independent and displacement as dependent, plot d vs t. Time Displacement 0. .5 1.0 1.5 2.0 2.5 3.0 3.5 1.2 4.9 11.0 19.5 30.7 44.3 59.8 9. Plot velocity versus time. Calculate the slope of the line on the graph itself Time velocity 1.0 1.5 2.0 2.5 3.0 3.5 m/s 0 4.8 9.8 14.5 19.7 24.5 29.6 34.3 18
Answer #1

Ans.

1)

41.57s + 0.016s = 41.586s, since the number of the significant digits is 2 in 0.016 s, so we round the answer to two significant figures. The 3rd digits in the sum is greater than 5, we round up the sum.

The sum = 42. s

----------------------------

2)

1.57 m x 12.02 m = 18.8714 m2 . SInce the number of the significant figures in 1.57 m is three we round the answer to three significant figures. The 4th digit in the product is greater than 5 so we round up the answer.

The product = 18.9 m2

-------------------------------------

3)

1502 kg - 3.14 kg = 1498.86 kg, The number of the significant figures in 3.14kg is three we round the answer to three significant figures. Writing the answer in scientific form: 1.49886*103 kg.

The 4th digit in the product is greater than 5 so we round the answer up.

The difference = 1.50x103 kg

--------------------------------

4)

Performing each division before adding them.

i) 999.0m/6.00s = 166.5 m/s and this rounds to 166 m/s (6.00s has 3 sig fig and the digit before 5 is even so we ignore 5)

ii) 888m/5.00s = 177.6 m/s and this rounds to 178 m/s

iii) 444m/4.80s = 92.5 m/s

Now we add them up

The answer = 166 + 178 + 92.5 = 436.5 m/s and this rounds to 436 m/s(answer)